# Bootstrapping Classical Shadows for Neural Quantum State Tomography

### v2

@@ I Introduction
Maximum likelihood estimation offers a path to alleviating all these challenges using memory-efficient parameterized models (i.e., ansatze). Model parameters are variationally optimized in update directions that make the training data most likely to be generated by the learned ansatze. Such variational ansatze can flexibly be used to impose physicality on the reconstructed states. Among various approaches, tensor networks and neural networks provide flexible architectures capable of capturing relevant regions within the Hilbert space. Moreover, when these networks possess an autoregressive property, they can efficiently provide independent samples for subsequent tasks downstream as generative models for the learned quantum states.
Recently, Ref. [2] proposed an approach for training a neural quantum state using classical shadows to combine the advantages of variational ansatze and classical shadow measurements. The authors introduce the infidelity between the classical shadow state (introduced in Ref. [1]) and the ansatz as their loss function, enabling the training of the ansatz on Clifford measurements. They suggest that, unlike the traditionally used cross-entropy loss function, infidelity offers a better optimization landscape due to its metric properties and bounded range. However, in the case of an unphysical state like the classical shadow state, the infidelity is not constrained within the physically valid $[0,1]$ range. Additionally, infidelity cannot guarantee bounded errors in approximating many quantities of interest from the trained model. Finally, the loss does not generalize to ansatze for mixed states.
Moreover, Ref. [2] faces important practical challenges in its training protocol. The trainability of the authors’ ansatz relies on an ad hoc initialization step. That is, they pretrain their model using the cross-entropy loss function on the classical dataset with measurements only in the computational basis, which was specifically chosen so for the GHZ state. This pretraining requires an additional set of measurements beyond the one used in the classical shadow technique. Furthermore, its success depends on the informational content of the underlying quantum state in the computational basis, rendering it non-universal.
@@ II.1 Neural Quantum States
is particularly appealing, as the resulting density matrix is physical. Here, $\bar{s}$ are auxiliary input variables representing the extended dimensions of the Hilbert space of the purified state. Therefore, their number $n_{\bar{s}}$ should reflect the expectation of mixedness of the target density matrix, with $n_{s}$ being a theoretically proven sufficient upper bound for capturing any density matrix over the physical input variables $s$. We treat $n_{\bar{s}}$ as a hyperparameter on the same footing as other parameters affecting the neural network’s architecture and hence its expressive power.
@@ II.2 Classical Shadows
The classical shadow $\mathcal{S}_{\rho}$ suffices to predict $M$ arbitrary linear and nonlinear functions of the state $\rho$ up to an additive error $\epsilon$ if it is of size $\Omega(\log(M)~{}{\rm max}_{i}~{}||O_{i}||^{2}_{\rm shadow}/\epsilon^{2})$. The shadow norm depends on the ensemble from which the unitaries $U_{i}$ are drawn for creating the classical shadow [1] and has efficient bounds, particularly if the ensemble satisfies a unitary 3-design property. For example, this is the case if the unitaries are drawn from the Clifford set as in Ref. [1] or using analog evolutions of different lengths of time as in Ref. [14].
Two particularly interesting ensembles are the Pauli and Clifford groups. These ensembles are computationally favourable because of their efficient classical simulations, according to the Gottesman–Knill theorem. The inverse twirling maps for the $n$-qubit Clifford and Pauli ensembles take the forms $\mathcal{M}^{-1}_{n}\left(X\right)=(2^{n}+1)X-\mathbb{I}$ and $\otimes_{k=1}^{n}\mathcal{M}_{1}^{-1}$, respectively. The shadow norm scales well for the Clifford ensemble, while the Pauli ensemble is more experimentally friendly. For the Pauli ensemble, the shadow norm scales exponentially poorly with respect to the locality of an observable, rendering the procedure inefficient in estimating high-weight observables. A middle-ground solution suggested in the literature [4, 5, 6] is to use shallow subsets of the Clifford group.
@@ II.3.1 Shadow-Based Loss Functions
The same loss function can also be modified for the case of Pauli ensembles as follows. For each Pauli $U_{i}=U_{i,1}\otimes\cdots\otimes U_{i,n}$ and any bitstring $b\in B_{n}\equiv\{0,1\}^{n}$, let $|b|=\sum_{i=1}^{n}b_{i}$ be the number of ones in $b$. Let $U_{i,b}=U_{i,1}^{b_{1}}\otimes\cdots\otimes U_{i,n}^{b_{n}}$ be the operator that has an identity factor for each zero index in $b$. In this case,
@@
We suppress the superscripts “$\rm{P}$” and “$\rm{C}$” in the rest of this section, as the discussion that follows applies to both training datasets. Unfortunately, these loss functions are not generalizable to mixed states. To circumvent this drawback, we reconsider the Kullback–Leibler (KL) divergence. Our Pauli or Clifford datasets can be viewed as samples drawn from a target probability measure $p^{\rm{data}}$:${}\rm{Stab}\to\mathbb{R}$ on the space of stabilizer states. For Pauli measurements this is the product of single-qubit stabilizer states $\rm{Stab}_{1}^{\otimes n}$, and for Clifford measurements it is $\rm{Stab}_{n}$, the full set of all $n$-qubit stabilizer states. The neural network (or, more generally, the autoregressive model) representing the quantum state is a priori a generative model with explicit access to a probability distribution $p_{\lambda}$:{${}\{0,1\}^{n}\to\mathbb{R}$ on the computational basis states. However, the efficiency of stabilizer state tableau representation allows us to extend it to a measure $p_{\lambda}$:${}\rm{Stab}\to\mathbb{R}$ on the stabilizer states. The KL divergence between $p_{\lambda}$ and $p^{\rm{data}}$ can be written as
@@
We note that the entropy of the underlying data distribution is a constant with respect to the variational parameters $\lambda$. Consequently, the loss function reduces to the cross-entropy term $-\langle\ln p_{\lambda}(\phi)\rangle_{\phi\sim p^{\mathrm{data}}}$.
Let $\mathcal{D}\subseteq\rm{Stab}$ be an independent identically distributed family of samples (allowing repetitions) drawn from $p^{\mathrm{data}}$, and $\widetilde{\mathcal{D}}$ be the set of elements in the dataset $\mathcal{D}$ (removing repetitions). The cross-entropy loss function can be approximated by the log-likelihood of $\mathcal{D}$,
@@ II.3.2 Monte Carlo Estimation and Gradients
Central to the computational tractability of the considered loss functions Eqs. 3, 5 and 8 is the evaluation of the overlap between a stabilizer state $\phi$ and the neural state $\psi_{\lambda}$:
@@
$\innerproduct{\psi_{\lambda}}{\phi}$ $=\expectationvalue{\frac{\nabla_{\lambda}\psi_{\lambda}^{*}(s)}{\phi^{*}(s)}}_{s\sim p_{\phi}}.$
@@ III.1 Scaling and Variance Study
The advantage of shadow weights in terms of estimating $p^{\rm{data}}_{\rm{emp}}$ is evident for smaller systems, as shown in Fig. 1 (a), though it gradually diminishes as the systems grow larger. In contrast, Fig. 1(b) indicates that this benefit can be reclaimed by increasing the number of measurements performed. A crucial factor in interpreting the plots in Fig. 1 is considering the difference in the configuration space dimensions of $p^{\rm{data}}_{\rm{emp}}$ associated with Pauli and Clifford measurements. Notably, Pauli measurements can reconstruct $4^{n}$ distinct stabilizers, whereas Clifford measurements can reconstruct $2^{0.5+\mathcal{O}(1)n^{2}}$ distinct stabilizers [aaronson2004improved]. For $n=6$ qubits, the dataset sizes considered in Fig. 1(b) must be compared to the dimension of the stabilizer space. For the Pauli distirbution, the dataset ranges from 5% to 500% of the total set size of distinct stabilizers. However, for the Clifford distribution, the maximum number of measurements is less than $0.001\%$ of the corresponding stabilizer set size. This significant difference explains the slight scaling advantage of the $p^{\rm{emp}}$ Pauli measurements over the $p^{\rm{emp}}$ Clifford measurements, as the former have a higher likelihood of obtaining repeated stabilizers, while the latter degenerate to a uniform distribution. For shadow weights, we observe a scaling advantage over empirical distributions in both regimes. However, unlike in the case of the empirical distributions, the Clifford distribution shows a slight improvement over the Pauli distribution. Such an advantage of Clifford shadow weights over Pauli ones, despite being in a regime with less data, is consistent with the theoretical expectation of an exponential benefit in predicting high-weight, low-rank observables (i.e., other stabilizers) [1]. Therefore, given a sufficiently large measurement dataset, training with shadow weights-based cross-entropy is expected to produce neural quantum states that better capture the underlying quantum state.
@@
We now evaluate the advantages of the proposed stabilizer-based method of Section II.3.2, over the neural state-based method of Eq. 10, for Monte Carlo estimation of the overlap, which is used to compute the training gradient. For benchmarking purposes, we analyze the errors in the gradient estimates relative to the true gradient, obtained through an exhaustive evaluation of the inner product Eq. 9. Specifically, we focus on the error in the gradient direction and distinguish between errors calculated over a mini-batch and a full batch of measurements. Figure 2 illustrates the evolution of these errors as the neural quantum state is trained using the true gradient. Figure 2(a) shows that both estimators struggle most with capturing the correct gradient direction during the early stages of training, where the neural quantum state is likely the most different from the target state. However, after several epochs, as the neural state becomes more similar to the target, the stabilizer-based estimation significantly improves, while the neural state estimation remains relatively high and inconsistent. Throughout the entire training process, the stabilizer-based estimator consistently provides estimates with much lower variance, which is essential for training stability. Additionally, our proposed estimator is consistently more accurate than the neural-based estimator. Being an unbiased estimator, the stochastic error due to mini-batching is expected to be cancelled out when averaged over the mini-batches and thus can be lowered using smaller learning rates, albeit at the cost of slower training dynamics. Figure 2(b) shows whether this expected behaviour holds for the given estimators. There is a clear difference between them: the stabilizer-based estimator functions as an unbiased estimator, reducing gradient errors by averaging over mini-batches. In contrast, the neural state-based estimator exhibits a systematic bias throughout the entire training duration, resulting in an extreme accumulation of errors from the true gradient. To eliminate this finite-sample bias, a significant increase in the samples generated by the neural state would be necessary. Therefore, given a fixed computational resource budget, our benchmarking results demonstrate that the stabilizer-based gradient estimation has a significant advantage over the neural state-based gradient estimation, highlighting its potential for optimization.
@@ III.2 Training Performance on Pure GHZ States
Figure 3: Benchmarking results of the performance of four methods in reconstructing both six- and eight-qubit GHZ states. We conduct our experiments using 1000 classical shadows based on Pauli or Clifford measurements (as specified in the titles above the plots), averaging the results over five training trials on the same dataset trained over 50 epochs. Each estimate of the overlap between a neural quantum state and a stabilizer (see Eq. 9) is based on 500 newly generated samples. The final states for each loss function are represented by markers. Triangle and square markers represent the samples from the last epoch for the stabilizer- and neural state-based infidelity loss functions, respectively. Circular and “x” markers represent the samples from the last epoch for the empirical and shadow based cross-entropy loss functions, respectively. The upper row displays the loss functions while the lower row shows the associated infidelity with respect to the true state as a measure of the generalization power of the neural quantum state.
@@
We now discuss the Clifford measurements, beginning with the benchmarking performed on the six-qubit GHZ state shown in Fig. 3(c) and (d). Our previous observations regarding the advantages of stabilizer-based sampling generalize to this scenario and, in fact, are applicable to the infidelity loss function as well. Indeed, the loss function $\mathcal{L}_{\text{inf}}^{\text{SS}}$ leads to an improved generalization with the final infidelity of roughly an order of magnitude compared to the loss function $\mathcal{L}_{\text{inf}}^{\text{NS}}$, as shown in Fig. 3(b). Interestingly, this improvement in generalization is achieved despite their corresponding loss functions yielding similar values at the end of training. Further examination of Fig. 3(c) and (d) reveals that our loss function $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ greatly outperforms the previously proposed shadow-based loss function $\mathcal{L}_{\text{inf}}^{\text{NS}}$ [2] in both the convergence rate and generalization. In addition, a comparison between the loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ reveals that our cross-entropy estimator leads to a superior convergence rate and lower infidelity, illustrating the benefits of shadow weights extracted from the classical shadow state (see Eq. (7)). Overall, both loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{inf}}^{\text{SS}}$ demonstrate the best convergence rates. The loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ exhibit the greatest generalization, approaching an infidelity of $0.01$, while our estimator shows the greatest stability in reaching states in proximity to the true states. An even more drastic difference in performance is observed for the eight-qubit GHZ state, shown in Fig. 3(e) and (f). In this case, the loss functions $\mathcal{L}_{\text{inf}}^{\text{NS}}$, $\mathcal{L}_{\text{inf}}^{\text{SS}}$, and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ struggle to escape local minima. While all of the loss functions result in infidelities that remain stuck at $0.99$, only the loss function $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ stands out, achieving an infidelity as low as $0.05$. To summarize, our results indicate that our proposed method provides the best training objective for neural networks trained to approximate an underlying quantum state from limited measurements.

### v3

@@ I Introduction
Maximum likelihood estimation offers a path to alleviating all these challenges using memory-efficient parameterized models (i.e., ansatze) and randomized subsets of tomographically complete datasets. Model parameters are variationally optimized in update directions that make the training data the most likely to be generated by the learned ansatze. Such variational ansatze can flexibly be used to impose physicality on the reconstructed states. Among various approaches, tensor networks and neural networks provide flexible architectures capable of capturing relevant regions within the Hilbert space. Moreover, when these networks possess an autoregressive property, they can efficiently provide independent samples for subsequent tasks downstream as generative models for the learned quantum states.
Recently, Ref. [2] proposed an approach for training a neural quantum state using classical shadows to combine the advantages of variational ansatze and classical shadow measurements. The authors introduce the infidelity between the classical shadow state (introduced in Ref. [1]) and the ansatz as their loss function, since it is provably efficient to estimate this quantity using Clifford shadow measurements. However, in the case of an unphysical state like the classical shadow state, the infidelity estimator is not constrained within the physically valid $[0,1]$ range. Additionally, infidelity cannot guarantee bounded errors in approximating many quantities of interest from the trained model. Finally, the loss does not generalize to ansatze for mixed states.
Moreover, the method in Ref. [2] faces important practical challenges in its training protocol. The trainability of the authors’ ansatz relies on an ad hoc initialization step. That is, they pretrain their model using the cross-entropy loss function on the classical dataset with measurements only in the computational basis, which works well for quantum states that are sparse in the computational basis, such as the GHZ state. This pretraining requires an additional set of measurements beyond the one used in the classical shadow technique. Furthermore, its success depends on the informational content of the underlying quantum state in the computational basis, rendering it non-universal.
@@ II.1 Neural Quantum States
is particularly appealing, as the resulting density matrix is physical. Here, $\bar{s}$ are auxiliary input variables representing the extended dimensions of the Hilbert space of the purified state. Therefore, their number $n_{\bar{s}}$ constrains the entropy of $\rho_{\lambda}$, with $n_{s}$ being a theoretically proven sufficient upper bound for capturing any density matrix over the physical input variables $s$. We treat $n_{\bar{s}}$ as a hyperparameter on the same footing as other parameters affecting the neural network’s architecture and hence its expressive power.
@@ II.2 Classical Shadows
The classical shadow $\mathcal{S}_{\rho}$ suffices to predict $M$ arbitrary linear and nonlinear functions $O_{i}$ of the state $\rho$ up to an additive error $\epsilon$ if it is of size $\Omega(\log(M)~{}{\rm max}_{i}~{}||O_{i}||^{2}_{\rm shadow}/\epsilon^{2})$. The shadow norm depends on the ensemble from which the unitaries $U_{i}$ are drawn for creating the classical shadow [1] and has efficient bounds, particularly if the ensemble satisfies a unitary 3-design property. For example, this is the case if the unitaries are drawn from the Clifford set as in Ref. [1] or using analog evolutions of different lengths of time as in Ref. [14].
Two particularly interesting ensembles are the Pauli and Clifford groups. These ensembles are computationally favourable because of their efficient classical simulations, according to the Gottesman–Knill theorem. The inverse twirling maps for the $n$-qubit Clifford and Pauli groups take the forms $\mathcal{M}^{-1}_{n}\left(X\right)=(2^{n}+1)X-\mathbb{I}$ and $\otimes_{k=1}^{n}\mathcal{M}_{1}^{-1}$, respectively. The shadow norm scales well for the Clifford ensemble, while the Pauli ensemble is more experimentally friendly. For the Pauli ensemble, the shadow norm scales exponentially poorly with respect to the locality of an observable, rendering the procedure inefficient in estimating high-weight observables. A middle-ground solution suggested in the literature [4, 5, 6] is to use shallow subsets of the Clifford group.
@@ II.3.1 Shadow-Based Loss Functions
The same loss function can also be modified for the case of Pauli ensembles as follows. For each Pauli $U_{i}=U_{i,1}\otimes\cdots\otimes U_{i,n}$ and any bitstring $b\in\mathcal{B}_{n}\equiv\{0,1\}^{n}$, let $|b|=\sum_{i=1}^{n}b_{i}$ be the number of ones in $b$. Let $U_{i,b}=U_{i,1}^{b_{1}}\otimes\cdots\otimes U_{i,n}^{b_{n}}$ be the operator that has an identity factor for each zero index in $b$. In this case,
@@
We suppress the superscripts “$\rm{P}$” and “$\rm{C}$” in the rest of this section, as the discussion that follows applies to both training datasets. Unfortunately, these loss functions are not generalizable to mixed states. To circumvent this drawback, we reconsider the Kullback–Leibler (KL) divergence. Our Pauli or Clifford datasets can be viewed as samples drawn from a target probability measure $p^{\rm{data}}$:${}\rm{Stab}\to\mathbb{R}$ on the space of stabilizer states. For Pauli measurements this is the product of single-qubit stabilizer states $\mathrm{Stab}_{1}^{\otimes n}$, and for Clifford measurements it is $\mathrm{Stab}_{n}$, the full set of all $n$-qubit stabilizer states. The neural network (or, more generally, the autoregressive model) representing the quantum state is a priori a generative model with explicit access to a probability distribution $p_{\lambda}$:{${}\{0,1\}^{n}\to\mathbb{R}$ on the computational basis states. However, the efficiency of stabilizer state tableau representation allows us to extend it to a measure $p_{\lambda}$:${}\rm{Stab}\to\mathbb{R}$ on the stabilizer states. The KL divergence between $p_{\lambda}$ and $p^{\rm{data}}$ can be written as
@@
We note that the entropy of the underlying data distribution is constant with respect to the variational parameters $\lambda$. Consequently, the loss function reduces to the cross-entropy term $-\langle\ln p_{\lambda}(\phi)\rangle_{\phi\sim p^{\mathrm{data}}}$.
Let $\mathcal{D}\subseteq\rm{Stab}$ be an independent, identically distributed family of samples (allowing repetitions) drawn from $p^{\mathrm{data}}$, and $\widetilde{\mathcal{D}}$ be the set of elements in the dataset $\mathcal{D}$ (removing repetitions). The cross-entropy loss function can be approximated by the log-likelihood of $\mathcal{D}$,
@@ II.3.2 Monte Carlo Estimation and Gradients
Central to the computational tractability of the considered loss functions Eqs. 3, 5, and 8 is the evaluation of the overlap between a stabilizer state $\phi$ and the neural state $\psi_{\lambda}$:
@@
$\nabla_{\lambda}\innerproduct{\psi_{\lambda}}{\phi}$ $=\expectationvalue{\frac{\nabla_{\lambda}\psi_{\lambda}^{*}(s)}{\phi^{*}(s)}}_{s\sim p_{\phi}}.$
@@ III.1 Scaling and Variance Study
The advantage of shadow weights in terms of estimating $p^{\rm{data}}_{\rm{emp}}$ is evident for smaller systems, as shown in Fig. 1 (a), though it gradually diminishes as the systems grow larger. In contrast, Fig. 1(b) indicates that this benefit can be reclaimed by increasing the number of measurements performed. A crucial factor in interpreting the plots in Fig. 1 is considering the difference in the configuration space dimensions of $p^{\rm{data}}_{\rm{emp}}$ associated with Pauli and Clifford measurements. Notably, Pauli measurements can reconstruct $4^{n}$ distinct stabilizers, whereas Clifford measurements can reconstruct $2^{0.5+\mathcal{O}(1)n^{2}}$ distinct stabilizers [15]. For $n=6$ qubits, the dataset sizes considered in Fig. 1(b) must be compared to the dimension of the stabilizer space. For the Pauli distribution, the dataset ranges from 5% to 500% of the total set size of distinct stabilizers. However, for the Clifford distribution, the maximum number of measurements is less than $0.001\%$ of the corresponding stabilizer set size. This significant difference explains the slight scaling advantage of the $p^{\rm{emp}}$ Pauli measurements over the $p^{\rm{emp}}$ Clifford measurements, as the former have a higher likelihood of obtaining repeated stabilizers, while the latter degenerate to a uniform distribution. For shadow weights, we observe a scaling advantage over empirical distributions in both regimes. However, unlike in the case of the empirical distributions, the Clifford distribution shows a slight improvement over the Pauli distribution. Such an advantage of Clifford shadow weights over Pauli ones, despite being in a regime with less data, is consistent with the theoretical expectation of an exponential benefit in predicting high-weight, low-rank observables (i.e., other stabilizers) [1]. Therefore, given a sufficiently large measurement dataset, training with shadow weights-based cross-entropy is expected to produce neural quantum states that better capture the underlying quantum state.
@@
We now evaluate the advantages of the proposed stabilizer-based method of Section II.3.2, over the neural state-based method of Eq. 10, for Monte Carlo estimation of the overlap, which is used to compute the training gradient. For benchmarking purposes, we analyze the errors in the gradient estimates relative to the true gradient, obtained through an exhaustive evaluation of the inner product of Eq. 9. Specifically, we focus on the error in the gradient direction and distinguish between errors calculated over a mini-batch and a full batch of measurements. Figure 2 illustrates the evolution of these errors as the neural quantum state is trained using the true gradient. Figure 2(a) shows that both estimators struggle most with capturing the correct gradient direction during the early stages of training, where the neural quantum state is likely the most different from the target state. However, after several epochs, as the neural state becomes more similar to the target, the stabilizer-based estimation significantly improves, while the neural state estimation remains relatively high and inconsistent. Throughout the entire training process, the stabilizer-based estimator consistently provides estimates with much lower variance, which is essential for training stability. Additionally, our proposed estimator is consistently more accurate than the neural-based estimator. Being an unbiased estimator, the stochastic error due to mini-batching is expected to be cancelled out when averaged over the mini-batches and thus can be lowered using smaller learning rates, albeit at the cost of slower training dynamics. Figure 2(b) shows whether this expected behaviour holds for the given estimators. There is a clear difference between them: the stabilizer-based estimator functions as an unbiased estimator, reducing gradient errors by averaging over mini-batches. In contrast, the neural state-based estimator exhibits a systematic bias throughout the entire training duration, resulting in an extreme accumulation of errors from the true gradient. To eliminate this finite-sample bias, a significant increase in the samples generated by the neural state would be necessary. Therefore, given a fixed computational resource budget, our benchmarking results demonstrate that the stabilizer-based gradient estimation has a significant advantage over the neural state-based gradient estimation, highlighting its potential for optimization.
@@ III.2 Training Performance on Pure GHZ States
Figure 3: Benchmarking results of the performance of four methods in reconstructing both six- and eight-qubit GHZ states. We conduct our experiments using 1000 classical shadows based on Pauli or Clifford measurements (as specified in the titles above the plots), averaging the results over five training trials on the same dataset trained over 50 epochs. Each estimate of the overlap between a neural quantum state and a stabilizer (see Eq. 9) is based on 500 newly generated samples. The final states for each loss function are represented by markers. Triangles and square markers represent the samples from the last epoch for the stabilizer- and neural state-based infidelity loss functions, respectively. Circles and “x” markers represent the samples from the last epoch for the empirical and shadow-based cross-entropy loss functions, respectively. The upper row displays the infidelity loss (left y-axis) and the cross-entropy loss (right y-axis) for the loss functions while the lower row shows the associated infidelity with respect to the true state as a measure of the generalization power of the neural quantum state. In order to facilitate comparison with the infidelity loss, the range of the cross-entropy loss is normalized not to exceed 1.
@@
We now discuss the Clifford measurements, beginning with the benchmarking performed on the six-qubit GHZ state shown in Fig. 3(c) and (d). Our previous observations regarding the advantages of stabilizer-based sampling generalize to this scenario and, in fact, are applicable to the infidelity loss function as well. Indeed, the loss function $\mathcal{L}_{\text{inf}}^{\text{SS}}$ leads to an improved generalization with the final infidelity of roughly an order of magnitude compared to the loss function $\mathcal{L}_{\text{inf}}^{\text{NS}}$, as shown in Fig. 3(d). Interestingly, this improvement in generalization is achieved despite their corresponding loss functions yielding similar values at the end of training. Further examination of Fig. 3(c) and (d) reveals that our loss function $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ greatly outperforms the previously proposed shadow-based loss function $\mathcal{L}_{\text{inf}}^{\text{NS}}$ [2] in both the convergence rate and generalization. In addition, a comparison between the loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ reveals that our cross-entropy estimator leads to a superior convergence rate and lower infidelity, illustrating the benefits of shadow weights extracted from the classical shadow state (see Eq. (7)). Overall, both loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{inf}}^{\text{SS}}$ demonstrate the best convergence rates. The loss functions $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ exhibit the greatest generalization, approaching an infidelity of $0.01$, while our estimator shows the greatest stability in reaching states in proximity to the true states. An even more drastic difference in performance is observed for the eight-qubit GHZ state, shown in Fig. 3(e) and (f). In this case, the loss functions $\mathcal{L}_{\text{inf}}^{\text{NS}}$, $\mathcal{L}_{\text{inf}}^{\text{SS}}$, and $\mathcal{L}_{\text{ECE}}^{\text{SS}}$ struggle to escape local minima. While all of the loss functions result in infidelities that remain stuck at $0.99$, only the loss function $\mathcal{L}_{\text{SCE}}^{\text{SS}}$ stands out, achieving an infidelity as low as $0.05$. To summarize, our results indicate that our proposed method provides the best training objective for neural networks trained to approximate an underlying quantum state from limited measurements.